### v2

# On the Theoretical Foundation of Sparse Dictionary Learning in Mechanistic Interpretability

Abstract
As AI models achieve remarkable capabilities across diverse domains, understanding what representations they learn and how they process information has become increasingly important for both scientific progress and trustworthy deployment. Recent works in mechanistic interpretability have shown that neural networks represent meaningful concepts as directions in their representation spaces and often encode diverse concepts in superposition. Various sparse dictionary learning (SDL) methods, including sparse autoencoders, transcoders, and crosscoders, address this by training auxiliary models with sparsity constraints to disentangle these superposed concepts into monosemantic features. These methods have demonstrated remarkable empirical success but have limited theoretical understanding. Existing theoretical work is limited to sparse autoencoders with tied-weight constraints, leaving the broader family of SDL methods without formal grounding. In this work, we develop the first unified theoretical framework considering SDL as one optimization problem. We demonstrate how diverse methods instantiate the theoretical framework and provide rigorous analysis of the optimization landscape. We provide novel theoretical explanations for empirically observed phenomena, including feature absorption and dead neurons. We design the Linear Representation Bench, a benchmark that strictly follows the Linear Representation Hypothesis, to evaluate SDL methods with fully accessible ground-truth features. Motivated by our theory and findings, we develop feature achoring, a novel technique applicable for all SDL methods, to enhance their feature recovery capabilities.
Sparse Autoencoders, Dictionary Learning, Interpretability, Neural Networks
1 Introduction
As artificial intelligence is increasingly deployed in high-stakes applications, from medical diagnosis to autonomous systems, the need for model interpretability has grown from an academic curiosity to a practical necessity (Lipton, 2017; Rudin, 2019). Numerous approaches have been proposed to peer inside the “black box,” including probing classifiers (Belinkov, 2021), attention visualization (Vaswani et al., 2023), concept bottleneck models (Koh et al., 2020), and feature attribution methods (Sundararajan et al., 2017). Among the efforts to enhance model interpretability, one line of research focuses on the structure of the model representations, pointing out that AI models often encode interpretable concepts as linear directions in superposition (Park et al., 2024; Elhage et al., 2022). As a consequence, individual neurons are typically polysemantic, responding to multiple unrelated concepts unlike their biological counterparts (Bricken et al., 2023; Templeton et al., 2024; Quiroga et al., 2005).
To address the polysemantic model representations, Sparse Dictionary Learning (SDL) has emerged as a promising paradigm (Sharkey et al., 2025). By training auxiliary models, such as sparse autoencoders (SAEs) (Cunningham et al., 2023), transcoders (Dunefsky et al., 2024), and crosscoders (Lindsey et al., 2024), that reconstruct neural activations through a sparsity-constrained bottleneck, SDL methods aim to identify the interpretable concepts underlying the polysemantic representations. These methods have achieved remarkable empirical success, scaling to frontier language models (Templeton et al., 2024; Gao et al., 2024) and enabling applications ranging from feature steering (Wang et al., 2025) to circuit analysis (O’Neill and Bui, 2024) and interpretable medical diagnosis (Abdulaal et al., 2024).
Despite these empirical advances, due to a lack of unified theoretical framework to analyze SDL methods, the field identifies various phenomena without rigorous theoretical explanation, including feature splitting and feature absorption (Chanin et al., 2025), the existence of dead neurons and the effectiveness of resampling (Bricken et al., 2023). Without theoretical grounding, the development of improved SDL methods remains largely empirical, potentially leaving highly effective techniques underexplored. While previous works on sparse coding and dictionary learning have established solid theoretical results (Spielman et al., 2012; Gribonval and Schnass, 2010), these results do not directly address SDL methods used in mechanistic interpretability. Recent work by Cui et al. (2025) has taken a key step by establishing necessary and sufficient conditions for identifiable sparse autoencoders under the linear representation hypothesis. However, their framework focuses exclusively on SAEs with tied-weight constraints, leaving other SDL methods without theoretical grounding.
In this work, we develop a unified theoretical framework that formalizes SDL as a general optimization problem, encompassing various SDL methods (Bussmann et al., 2024, 2025; Tang et al., 2025b) as special cases. We demonstrate how these diverse methods instantiate our framework through different choices of input-output representation pairs, activation functions, and loss designs. We establish rigorous conditions under which SDL methods provably recover ground-truth interpretable features, characterizing the roles of feature sparsity, latent dimensionality, and activation functions. Through detailed analysis of the optimization landscape, we demonstrate that global minima correspond to correct feature recovery while spurious local minima exhibit un-disentangled polysemanticity, providing novel theoretical explanation for feature absorption (Chanin et al., 2025) and neuron resampling (Bricken et al., 2023). We design the Linear Representation Bench, a synthetic benchmark that strictly follows the Linear Representation Hypothesis, to evaluate SDL methods with fully accessible ground-truth features. Motivated by our theoretical insights, we propose feature anchoring that achieve improved feature recovery by constraining learned features to known anchor directions. Our contributions are as follows:
  • We build the first theoretical framework for SDL in mechanistic interpretability as a general optimization problem encompassing diverse SDL methods.
  • We provide rigorous theoretical results on the optimization landscape and provide novel explanations for various phenomena observed empirically.
  • We design the Linear Representation Bench, a benchmark enabling controlled evaluation of SDL methods with fully accessible ground-truth features.
  • We propose feature anchoring that constrains learned features to known anchor directions, achieving improved feature recovery performance.
2 Preliminaries
In this section, we present a unified theoretical framework for Sparse Dictionary Learning (SDL). We begin with the formal definitions of foundational concepts and the linear representation hypothesis. Then we introduce our framework and how various SDL methods instantiate it.
2.1 Input Distribution and Model Representation
Definition 2.1 (Input Distribution).
Let $\mathcal{D}$ denote the distribution over possible inputs to a neural network. For example, $\mathcal{D}$ could be the distribution of natural images or the distribution of text sequences.
Definition 2.2 (Model Representation).
For a given model representation $\mathbf{x}$, let $n\in\mathbb{N}$ be its dimensionality. For each $s\sim\mathcal{D}$, the network produces a representation vector $\mathbf{x}(s)$, which is directly observable by running the model on $s$.
2.2 The Linear Representation Hypothesis
Empirical studies in mechanistic interpretability have observed that neural network representations encode meaningful concepts as linear directions, often in superposition (Marks and Tegmark, 2024; Nanda et al., 2023; Jiang et al., 2024; Park et al., 2025). Following Elhage et al. (2022) and Park et al. (2024), we formalize the hypothesis as follows.
Assumption 2.3 (Linear Representation Hypothesis).
A model representation $\mathbf{x}_{p}\in\mathbb{R}^{n_{p}}$ satisfies the Linear Representation Hypothesis (LRH) if there exists a feature function $\mathbf{x}:\mathcal{D}\to\mathbb{R}^{n}$ and a feature matrix $W_{p}\in\mathbb{R}^{n_{p}\times n}$ such that:
  1. 1.
    Linear Decomposition: For all $s\sim\mathcal{D}$,
    $\mathbf{x}_{p}(s)=W_{p}\mathbf{x}(s).$
  2. 2.
    Non-negativity: $\mathbf{x}(s)\geq 0$ component-wise.
  3. 3.
    Sparsity: There exists $S\in[0,1]$ such that $\forall i\in[n]$,
    $\Pr_{s\sim\mathcal{D}}(x_{i}(s)=0)\geq S.$
  4. 4.
    Interpretability: Each component $x_{i}$ of $\mathbf{x}$ corresponds to a human-interpretable concept.
We refer to $\mathbf{x}(s)$ as the ground-truth features, where each $x_{i}(s)$ represents the activation level of concept $i$ for input $s$. $W_{p}$ encodes how these features are represented in $\mathbf{x}_{p}$.
Assumption 2.4 (Representation Assumptions).
Following recent works (Elhage et al., 2022; Park et al., 2024), we assume the following conditions on the model representations satisfying the Linear Representation Hypothesis, which align with practical use cases:
  1. 1.
    Unit Norm: The feature matrix $W_{p}\in\mathbb{R}^{n_{p}\times n}$ has unit-norm columns:
    $\|W_{p}[:,i]\|_{2}=1\quad\forall i\in[n]$
  2. 2.
    Bounded Interference: Define the maximum interference as:
    $M:=\max_{i\neq j}\langle W_{p}[:,i],W_{p}[:,j]\rangle$
    In typical superposition scenarios, $M>0$ but remains small (close to zero), characterizing a regime where many interpretable features are compressed into fewer dimensions with bounded mutual interference.
  3. 3.
    Extreme Sparsity: The sparsity level $S$ approaches 1:
    $S\to 1$
  4. 4.
    Independence: The ground-truth features satisfy: For any input $s\sim\mathcal{D}$, the features $\{x_{i}(s)\}_{i=1}^{n}$ are mutually independent and independent of the sparsity level $S$.
where $W_{p}[:,i]$ denotes the $i$-th column of $W_{p}$.
2.3 General Optimization Framework for SDL
Now we formalize SDL as an optimization problem under the Linear Representation Hypothesis (Assumption 2.3) and the Representation Assumptions (Assumption 2.4).
Definition 2.5 (Sparse Dictionary Learning).
A SDL model maps an input representation $\mathbf{x}_{p}\in\mathbb{R}^{n_{p}}$ to a target representation $\mathbf{x}_{r}\in\mathbb{R}^{n_{r}}$ through a two-layer architecture:
  1. (i)
    An encoder layer that maps $\mathbf{x}_{p}$ to a latent space:
    $\mathbf{x}_{q}=\sigma(W_{e}\mathbf{x}_{p})$ (1)
  2. (ii)
    A decoder layer that maps the latents to $\mathbf{x}_{r}$:
    $\hat{\mathbf{x}}_{r}=W_{d}\mathbf{x}_{q}$ (2)
where $\mathbf{x}_{q}\in\mathbb{R}^{n_{q}}$, $W_{e}\in\mathbb{R}^{n_{q}\times n_{p}}$, $W_{d}\in\mathbb{R}^{n_{r}\times n_{q}}$, and $\sigma:\mathbb{R}^{n_{q}}\to\mathbb{R}^{n_{q}}$ is a sparsity-inducing activation function.
The SDL objective minimizes expected reconstruction error:
$\mathcal{L}_{\text{SDL}}=\mathbb{E}_{s\sim\mathcal{D}}\left[\|\mathbf{x}_{r}(s)-W_{d}\sigma(W_{e}\mathbf{x}_{p}(s))\|_{2}^{2}\right]$ (3)
Under the Linear Representation Hypothesis (Assumption 2.3), both representations admit linear decompositions $\mathbf{x}_{p}(s)=W_{p}\mathbf{x}(s)$ and $\mathbf{x}_{r}(s)=W_{r}\mathbf{x}(s)$ in terms of ground-truth features $\mathbf{x}(s)$. The loss can thus be expressed as:
$\mathcal{L}_{\text{SDL}}=\mathbb{E}_{s\sim\mathcal{D}}\left[\|W_{r}\mathbf{x}(s)-W_{d}\sigma(W_{e}W_{p}\mathbf{x}(s))\|_{2}^{2}\right]$ (4)
2.4 Instantiations: Existing SDL Methods
We now demonstrate how existing SDL methods instantiate our framework through adopting different choices of input-target pairs $(\mathbf{x}_{p},\mathbf{x}_{r})$ and activation functions $\sigma$, and proposing variants on the loss function $\mathcal{L}_{\text{SDL}}$.
Sparse Autoencoders (SAEs). SAEs (Cunningham et al., 2023) decompose polysemantic activations into monosemantic components through sparsity constraints. In our framework, SAEs are characterized by setting $\mathbf{x}_{r}=\mathbf{x}_{p}$ (self-reconstruction). The encoder projects to a higher-dimensional sparse latent space, encouraging $\mathbf{x}_{q}(s)$ to capture the underlying ground-truth features $\mathbf{x}(s)$ (Figure 1).
Figure 1: Sparse Autoencoder: encoder $W_{e}$ maps $\mathbf{x}_{p}$ to sparse latents $\mathbf{x}_{q}$, decoder $W_{d}$ reconstructs from $\mathbf{x}_{q}$.
Transcoders. Transcoders (Dunefsky et al., 2024; Paulo et al., 2025) capture interpretable features in layer-to-layer transformations. Unlike SAEs, transcoders approximate the inputoutput function of a target component, such as an an MLP, using a sparse bottleneck. In our proposed theoretical framework, transcoders set $\mathbf{x}_{p}=\text{NN}_{\ell}(s)$ and $\mathbf{x}_{r}=\text{NN}_{\ell+1}(s)$, where $\text{NN}_{\ell}$ denotes the target component’s input, and $\text{NN}_{\ell+1}$ denotes its output (Figure 2).
Figure 2: Transcoder: encoder $W_{e}$ maps $\text{NN}_{\ell}(s)$ to sparse latents $\mathbf{x}_{q}$, decoder $W_{d}$ predicts $\text{NN}_{\ell+1}(s)$.
Crosscoders. Crosscoders (Lindsey et al., 2024) discover shared features across multiple representation sources by jointly encoding and reconstructing concatenated representations. In our framework, crosscoders set $\mathbf{x}_{p}=[\mathbf{x}_{p}^{(1)};\ldots;\mathbf{x}_{p}^{(m)}]$ and $\mathbf{x}_{r}=[\mathbf{x}_{r}^{(1)};\ldots;\mathbf{x}_{r}^{(m)}]$ where each superscript denotes a different source (Figure 3).
Figure 3: Crosscoder: encoder $W_{e}$ maps concatenated multi-layer input $\mathbf{x}_{p}$ to $\mathbf{x}_{q}$, decoder $W_{d}$ reconstructs multi-layer output $\mathbf{x}_{r}$.
Variants of SDL Methods. Various SDL methods fit into our theoretical framework but differ in their choices of activation functions and loss designs. Bricken et al. (2023) and Templeton et al. (2024) use ReLU activation $\sigma_{\text{ReLU}}(z)=\max(0,z)$ with $L_{1}$ regularization on latents:
$\mathcal{L}=\mathbb{E}_{s\sim\mathcal{D}}\left[\|\mathbf{x}_{r}(s)-W_{d}\sigma_{\text{ReLU}}(W_{e}\mathbf{x}_{p}(s))\|_{2}^{2}+\lambda\|\mathbf{x}_{q}(s)\|_{1}\right]$ (5)
Rajamanoharan et al. (2024b) propose to utilize JumpReLU as the activation function:
$\sigma_{\text{JumpReLU}}(z)=z\cdot H(z-\theta)$ (6)
combined with a smoothed Heaviside function to directly penalize the $L_{0}$ norm. Makhzani and Frey (2014) introduce Top-$k$ activation:
$\sigma_{\text{Top-}k}(z)_{i}=\begin{cases}z_{i}&\text{if }z_{i}\text{ is among the }k\text{ largest components}\\ 0&\text{otherwise}\end{cases}$ (7)
Bussmann et al. (2024) extend this to Batch Top-$k$ activation, which applies Top-$k$ selection across a batch to allow different samples to have different numbers of activated features. Rajamanoharan et al. (2024a) use ReLU activation with an additional Heaviside gating function. Gao et al. (2024) employ Top-$k$ activation and introduce an auxiliary loss using Top-$k_{\text{aux}}$ dead latents:
$\mathcal{L}=\mathbb{E}_{s\sim\mathcal{D}}\Big[\|\mathbf{x}_{r}(s)-W_{d}\sigma_{\text{Top-}k}(W_{e}\mathbf{x}_{p}(s))\|_{2}^{2}$ (8)
$+\lambda_{\text{aux}}\|\mathbf{x}_{p}(s)-W_{d}^{\prime}\sigma_{\text{Top-}k_{\text{aux}}}(W_{e}\mathbf{x}_{p}(s))\|_{2}^{2}\Big]$
to resurrect dead neurons. Bussmann et al. (2025) and Tang et al. (2025b) use multiple $k$ values with a multi-scale loss:
$\mathcal{L}=\sum_{i=1}^{m}\lambda_{i}\mathbb{E}_{s\sim\mathcal{D}}\left[\|\mathbf{x}_{r}(s)-W_{d}\sigma_{\text{Top-}k_{i}}(W_{e}\mathbf{x}_{p}(s))\|_{2}^{2}\right]$ (9)
that sums reconstruction errors of different sparsity levels.
The variants described above demonstrate that diverse SDL methods can be unified under our general framework (Definition 2.5) through specific choices of activation functions $\sigma$ and loss modifications. This motivates our theoretical approach: rather than analyzing each variant separately, we establish a key property for all these activation functions.
Key Activation Function Property. Our analysis focuses on activation functions satisfying:
$\sigma(z)_{i}\in\{0,z_{i}\}\quad\forall i\in[n_{q}]$ (10)
This property holds for ReLU, JumpReLU, Top-$k$, Batch Top-$k$, and their compositions, the primary sparsity mechanisms in practice. Notably, all these activation functions can be expressed in the form $\sigma_{\text{Jump}}(z;c)$ for some threshold $c\geq 0$:
$\sigma_{\text{Jump}}(z;c)_{i}=z_{i}\cdot\mathbb{1}(z_{i}>c)$ (11)
3 Theoretical Results
Despite the empirical success of SDL methods across diverse applications, a significant gap exists between their practical use and our theoretical understanding of their optimization dynamics. This gap has important consequences: practitioners employ techniques like dead neuron resampling (Bricken et al., 2023) and observe phenomena like feature absorption (Chanin et al., 2025) without rigorous explanations for why these occur or how to systematically address them. Without theoretical grounding, the development of improved SDL methods remains largely empirical, potentially leaving highly effective techniques underexplored.
Our theoretical analysis bridges this gap by characterizing the SDL optimization landscape under the Linear Representation Hypothesis. Section 3.1 derives the necessary and sufficient conditions for the global minima of the optimization problem. Section 3.2 demonstrates the existence of spurious local minima and how they relate with polysemanticity. Section 3.3 discusses about feature absorption and hierarchical concepts. We provide full proofs in Appendix.
3.1 Global Minimum
Here we analyze the global minima of our proposed framework. Theorem 3.1 shows that $\mathcal{L}_{\text{SDL}}$ is well-approximated by an alternative loss on feature reconstructions, $\tilde{\mathcal{L}}_{\text{SDL}}$.
Theorem 3.1 (Loss Approximation).
Under Assumption 2.3 and Assumption 2.4, define the approximate Loss, $\tilde{\mathcal{L}}_{\text{SDL}}$, as:
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{d},W_{e}):=\sum_{d=1}^{n}M_{d}\left\|w_{r}^{d}-W_{d}\sigma(W_{e}w_{p}^{d})\right\|^{2}$ (12)
where $M_{d}=\Pr(\mathbf{x}(s)=x_{d}(s)e_{d})\cdot\mathbb{E}[x_{d}(s)^{2}|x_{d}(s)>0]$. Then
$|\mathcal{L}_{\text{SDL}}-\tilde{\mathcal{L}}_{\text{SDL}}|\leq O((1-S)^{2})$ (13)
Theorem 3.2 then demonstrates that the configuration $(W_{d}^{*},W_{e}^{*})=([W_{r},\mathbf{0}],[W_{p}^{\top};\mathbf{0}])$ achieves approximate loss bounded by $O((1-S)M^{2})$, establishing it as a near-optimal solution that recovers the ground-truth feature structure.
Theorem 3.2 (Approximate Loss at Global Minimum Configuration).
Consider the approximate SDL loss
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{d},W_{e})=\sum_{d=1}^{n}M_{d}\left\|w_{r}^{d}-W_{d}\sigma(W_{e}w_{p}^{d})\right\|^{2}$ (14)
where $M_{d}=\Pr(\mathbf{x}(s)=x_{d}(s)e_{d})\cdot\mathbb{E}[x_{d}(s)^{2}|x_{d}(s)>0]$.
When $n_{q}\geq n$ and $\sigma$ satisfies $\sigma(z)_{i}\in\{0,z_{i}\}$ for all $i$, the configuration
$W_{d}^{*}=[W_{r},\mathbf{0}],\quad W_{e}^{*}=\begin{bmatrix}W_{p}^{\top}\\ \mathbf{0}\end{bmatrix}$ (15)
where $\mathbf{0}$ denotes zero padding to dimension $n_{q}$, satisfies:
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{d}^{*},W_{e}^{*})=O((1-S)M^{2})$ (16)
where $M$ is the maximum interference.
Notice that the necessary and sufficient conditions for $\tilde{\mathcal{L}}_{\text{SDL}}=0$ require $w_{r}^{d}=W_{d}\sigma(W_{e}w_{p}^{d})$ for all $d\in[n]$. This system of $n$ vector equations in the encoder-decoder parameters is underdetermined, admitting multiple solutions beyond $(W_{d}^{*},W_{e}^{*})$. Crucially, some solutions achieve zero reconstruction loss without recovering any ground-truth features, as demonstrated in Figure 4. This underdetermined nature motivates feature anchoring, a novel technique that applies additional constraints to guide optimization.
Figure 4: Zero reconstruction loss without recovering ground-truth features. We design the Linear Representation Bench that enable full knowledge of the ground truth features to study SDL methods. We observe one concerning phenomenon that these methods can achieve zero loss without recovering any ground truth features. Left: Four ground-truth feature directions. Middle: Learned encoder directions fail to align with ground truth. Right: Learned decoder directions are rotated accordingly. Although $\tilde{\mathcal{L}}_{\text{SDL}}\approx 0$, the learned features bear no correspondence to interpretable ground-truth concepts, demonstrating the underdetermined nature of SDL optimization.
3.2 Polysemanticity and Local Minima
We now establish that SDL optimization can converge to spurious local minima with loss strictly greater than the global optimum. We also point out the connection between local minima and un-disentangled polysemanticity.
Example 3.3 (Spurious Local Minimum).
Consider $n=n_{p}=n_{q}=n_{r}=2$, $\sigma=\sigma_{\text{Top-1}}\cdot\sigma_{\text{ReLU}}$, and extreme sparsity ($S\to 1$). Let $W_{e}=\begin{bmatrix}(w_{e}^{1})^{\top}\\ (w_{e}^{2})^{\top}\end{bmatrix}$ and $W_{d}=[w_{d}^{1},w_{d}^{2}]$. The approximate loss decomposes as:
$\begin{split}\tilde{\mathcal{L}}_{\text{SDL}}(W_{d},W_{e})=M_{1}\left\|w_{r}^{1}-W_{d}\sigma(W_{e}w_{p}^{1})\right\|^{2}\\ +M_{2}\left\|w_{r}^{2}-W_{d}\sigma(W_{e}w_{p}^{2})\right\|^{2}\end{split}$ (17)
where $M_{1},M_{2}>0$ are constants.
Consider a configuration where:
$\langle w_{e}^{1},w_{p}^{1}\rangle>\langle w_{e}^{2},w_{p}^{1}\rangle>0,\quad\langle w_{e}^{1},w_{p}^{2}\rangle>\langle w_{e}^{2},w_{p}^{2}\rangle>0$ (18)
Under this condition, neuron 1 activates for both features while neuron 2 remains dead. Let $a_{1}=\langle w_{e}^{1},w_{p}^{1}\rangle$ and $a_{2}=\langle w_{e}^{1},w_{p}^{2}\rangle$. The optimal decoder is:
$w_{d}^{1,*}=\frac{M_{1}a_{1}w_{r}^{1}+M_{2}a_{2}w_{r}^{2}}{M_{1}a_{1}^{2}+M_{2}a_{2}^{2}}$ (19)
Since $w_{r}^{1}\not\propto w_{r}^{2}$, the resulting loss is:
$\tilde{\mathcal{L}}_{\text{SDL}}^{*}=\frac{M_{1}M_{2}(M_{2}a_{2}^{2}+M_{1}a_{1}^{2})\|a_{2}w_{r}^{1}-a_{1}w_{r}^{2}\|^{2}}{(M_{1}a_{1}^{2}+M_{2}a_{2}^{2})^{2}}>0$ (20)
The spurious local minimum in Example 3.3 demonstrates a prevalent failure mode for SDL methods: a single SDL latent neuron becomes polysemantic, responding to multiple ground-truth features, while other neurons remain dead. This occurs when optimization becomes trapped in a local minimum where gradient descent cannot escape. We now formalize and characterize this polysemanticity.
Definition 3.4 (Polysemanticity Pattern).
A polysemanticity pattern is a collection $(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ where each $\mathcal{F}_{i}\subseteq[n]$ represents a subset of ground-truth features. The pattern is called polysemantic if $\exists i,|\mathcal{F}_{i}|\geq 2$.
Given an SDL configuration with activation function $\sigma$ and projection vectors $\{w_{p}^{d}\}_{d=1}^{n}$, a polysemanticity pattern $(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ forming a partition of $[n]$ is called realizable if there exists an encoder $W_{e}=[(w_{e}^{1})^{\top},\ldots,(w_{e}^{n_{q}})^{\top}]^{\top}$ such that:
$\forall i\in[n_{q}],\quad\mathcal{F}_{i}=\left\{d\in[n]:\left(\sigma(W_{e}w_{p}^{d})\right)_{i}>0\right\}$ (21)
That is, neuron $i$ activates for exactly the features in $\mathcal{F}_{i}$ when they are active in isolation.
Theorem 3.5 (Existence of Spurious Local Minima).
Consider $\tilde{\mathcal{L}}_{\text{SDL}}(W_{d},W_{e})=\sum_{d=1}^{n}M_{d}\left\|w_{r}^{d}-W_{d}\sigma(W_{e}w_{p}^{d})\right\|^{2}$ under Assumptions 2.3 and 2.4 with $n\geq 2$ and $n_{q}\geq n$.
For any realizable polysemanticity pattern $(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ with at least one $|\mathcal{F}_{i}|\geq 2$, there exists a local minimum point $(W_{d},W_{e})$ of the framework exhibiting this pattern.
3.3 Feature Absorption and Hierarchical Concepts
Empirical observations suggest that feature absorption is prevalent during the training of SDL (Chanin et al., 2025). We argue that this phenomenon arises when ground-truth features exhibit hierarchical structure and can be considered as a special type of realizable polysemanticity pattern.
Example 3.6 (Feature Absorption).
Consider a representation space containing animal concepts. Ground-truth features include a parent concept ”Dog” and four sub-concepts: ”Border Collie”, ”Golden Retriever”, ”Husky”, and ”German Shepherd”.
As illustrated in Figure 5, in the ideal case, SDL would learn four separate monosemantic neurons for Dog, Cat, Horse, and Elephant, with each dog breed activating only the Dog neuron. However, feature absorption can occur during training, resulting in the pattern shown in Figure 5 (right): one neuron exclusively captures ”Border Collie” (the absorbed feature), while a second neuron responds to the remaining three breeds collectively (the main line interpretation of ”Dog”).
Definition 3.7 (Hierarchical Concept Structure).
A set of ground-truth features exhibits hierarchical structure if it contains a $p$ and a set of sub-concepts $c_{1},\ldots,c_{k}$ satisfying: $p(x)>0\iff\exists i\in[k],c_{i}(x)>0$.
Theorem 3.8 (Existence of Feature Absorption).
Any ground truth features exhibiting hierarchical structure correspond to a polysemanticity pattern with feature absorption. If realizable, there is a local minimum point exhibiting it.
Theorem 3.5 reveals the prevalence of local minima in SDL training that challenge gradient-based optimization and their connection to polysemantic neurons. Theorem 3.8 shows that hierarchical concept structures naturally give rise to feature absorption patterns, which manifest as local minima when realizable, explaining its empirical prevalence.
Figure 5: Feature absorption emerges from hierarchical concept structure. Left: Ideal SDL features without absorption. Right: hierarchical concept structure exists and only a proportion of the sub-concepts of ”Dog” can activate the SDL feature.
4 The Feature Anchoring
Motivated by our theoretical results, we propose Anchored Sparse Autoencoders (See Figure 6) to guide optimization toward the global optimum. Specifically, we constrain the first $k$ encoder rows and decoder columns to match ground-truth feature directions via an auxiliary loss to mitigate the undetermined nature of the SDL optimization problem:
$\begin{split}\mathcal{L}_{\text{anchor}}&=\|\mathbf{W}_{e}[1{:}k,:]-\tilde{\mathbf{W}}_{p}[:,1{:}k]^{\top}\|^{2}_{F}\\ &+\|\mathbf{W}_{d}[:,1{:}k]-\tilde{\mathbf{W}}_{r}[:,1{:}k]\|^{2}_{F}\end{split}$ (22)
These anchor features are obtained either from ground-truth when available (Linear Representation Bench) or by computing mean embeddings of dataset subpopulations (Luo et al., 2024). During the evaluation, we only compute the recovery metric for features that were not selected as anchors.
Figure 6: Anchored SAE. First $k$ neurons (highlighted) constrained to ground-truth via $\mathcal{L}_{\text{anchor}}$; other neurons learn via SDL.
5 Experimental Results
5.1 The Linear Representation Bench
To validate our theoretical predictions, we design the Linear Representation Bench, a benchmark that precisely instantiates Assumptions 2.3 and 2.4. We generate synthetic representations $\mathbf{x}_{p}(s)=\mathbf{W}_{p}\mathbf{x}(s)$ where ground-truth features $\mathbf{x}(s)\in\mathbb{R}^{n}$ follow shifted exponential distributions with sparsity $\mathcal{S}$. The feature matrix $\mathbf{W}_{p}\in\mathbb{R}^{n_{p}\times n}$ is constructed via gradient-based optimization to satisfy bounded interference $|\langle\mathbf{w}^{i}_{p},\mathbf{w}^{j}_{p}\rangle|\leq\epsilon$ or negative interference $\langle\mathbf{w}^{i}_{p},\mathbf{w}^{j}_{p}\rangle\leq 0$ with unit-norm columns. We evaluate SDL methods by measuring feature recovery: $\mathbb{E}_{i}[\max_{j}|\langle\mathbf{w}^{i}_{p},\mathbf{w}^{j}_{e}\rangle|]$, the expected maximum alignment between each ground-truth feature $\mathbf{w}^{i}_{p}$ and learned encoder columns $\mathbf{w}^{j}_{e}$. As shown in Table 1, our proposed Anchored SAE significantly improves feature recovery performance.
Interference n Method
SAE Anchored SAE
Bounded 1000 0.256 0.284
Negative 1000 0.235 0.260
Table 1: We compare standard SAE and Anchored SAE (with $k=30$ anchor features) on synthetic representations in $d=768$ dimensions with $n\in\{1000,1500,2000\}$ features (Appendix C).
5.2 Neuron Resampling Helps Escape Local Minima
Our local minima analysis (Theorem 3.5) provides a novel explanation for why dead neuron resampling (Bricken et al., 2023) improves SDL training. Dead neurons ($\mathcal{F}_{i}=\emptyset$) contribute zero gradient in spurious local minima, leaving optimization trapped. Resampling reinitializes these neurons toward under-reconstructed directions, perturbing the optimization away from the local basin. To validate this, we conduct experiments on the Llama 3 model (Figure 7). Resampling enables the optimizer to escape spurious local minima, achieving lower reconstruction loss.
Figure 7: Feature resampling accelerates convergence and improves final loss. Training curves on Llama 3 comparing standard SAE training (blue) with periodic dead neuron resampling (red).
6 Related Works
6.1 Mechanistic Interpretability
Interpretability is crucial for deploying AI in high-stakes domains such as medical diagnosis and financial modeling, where understanding model decisions is essential for safety and trust (Simon and Zou, 2024; Abdulaal et al., 2024). Traditional approaches include interpretability-by-design methods like Concept Bottleneck Models (Koh et al., 2020) and decision trees (Mienye and Jere, 2024), and post-hoc explanation methods like GradCAM (Selvaraju et al., 2019) and SHAP (Lundberg and Lee, 2017). Mechanistic interpretability (Sharkey et al., 2025; Bereska and Gavves, 2024) aims to reverse-engineer neural networks by understanding their internal computational mechanisms. SAEs (Shu et al., 2025) and related dictionary learning methods (Tang et al., 2025a; Dunefsky et al., 2024) decompose neural activations into sparse, interpretable features. Circuit analysis (Olah et al., 2020; Olsson et al., 2022) investigates how these features compose into computational algorithms.
6.2 Sparse Dictionary Learning
Sparse dictionary learning has a rich history predating its application to mechanistic interpretability. K-SVD (Aharon et al., 2006) established foundational methods for learning overcomplete dictionaries, while theoretical work in compressed sensing (Donoho, 2006) characterized recovery conditions, with Spielman et al. (2012) providing polynomial-time algorithms for exact reconstruction under sparsity assumptions. Safran and Shamir (2018) demonstrated that spurious local minima are common even in simple two-layer ReLU networks, highlighting optimization challenges that persist in modern applications. Recent work has adapted these principles to mechanistic interpretability. SAEs (Cunningham et al., 2023) apply dictionary learning to language model activations. Various variants are further developed, including transcoders (Dunefsky et al., 2024), crosscoders (Gao et al., 2024), Matryoshka SAEs (Bussmann et al., 2025), and hybrid approaches like Language-Grounded Sparse Encoders (Tang et al., 2025a). These methods have found applications beyond language models, including protein structure analysis (Simon and Zou, 2024; Gujral et al., 2025), medical imaging (Abdulaal et al., 2024), model evaluation (Tang et al., 2025b), board game analysis (Karvonen et al., 2024), and fMRI data analysis Mao et al. (2025).
7 Conclusion
We propose the first unified theoretical framework for Sparse Dictionary Learning in mechanistic interpretability, and demonstrate how various SDL methods instantiate our framework. Under the Linear Representation Hypothesis with bounded feature interference, we established conditions for provable recovery of interpretable features and characterized the optimization landscape. We proved that spurious local minima correspond to feature absorption, providing the first theoretical explanation for the connection between feature absorption and local minima. To validate our theory, we designed the Linear Representation Bench, a benchmark that strictly instantiates our theoretical assumptions with fully accessible ground-truth features. Motivated by our theoretical findings and numerical experiments, we proposed feature anchoring, applicable for all SDL methods, and achieving improved feature recovery performance.
Appendix A Proof of Theorem 3.1
Proof.
We establish how the loss decomposes when $S\to 1$. Let $W_{r}=[w_{r}^{1},w_{r}^{2},\ldots,w_{r}^{n}]$ and $W_{p}=[w_{p}^{1},w_{p}^{2},\ldots,w_{p}^{n}]$ denote the column representations of the reconstruction and projection matrices.
Under extreme sparsity ($S\to 1$), we consider at most one component $x_{d}(s)$ is active for any given input, and consider other scenarios as of low probability. The expectation decomposes over activation patterns:
$\mathcal{L}_{\text{SDL}}$ $=\mathbb{E}_{s}\left\|W_{r}\mathbf{x}(s)-W_{d}\sigma(W_{e}W_{p}\mathbf{x}(s))\right\|^{2}$ (23)
$=\sum_{m=0}^{n}\Pr(\|\mathbf{x}(s)\|_{0}=m)\cdot\mathbb{E}[\|\mathbf{x}_{r}(s)-W_{d}\sigma(W_{e}\mathbf{x}_{p}(s))\|^{2}\mid\|\mathbf{x}(s)\|_{0}=m]$ (24)
$=\Pr(\|\mathbf{x}(s)\|_{0}=0)\cdot 0$ (25)
$\quad+\sum_{d=1}^{n}\Pr(\mathbf{x}(s)=x_{d}(s)\mathbf{e}_{d})\cdot\mathbb{E}_{x_{d}(s)}\left[x_{d}(s)^{2}\left\|w_{r}^{d}-W_{d}\sigma(W_{e}w_{p}^{d})\right\|^{2}\mid x_{d}(s)>0\right]$ (26)
$\quad+\sum_{m=2}^{n}\Pr(\|\mathbf{x}(s)\|_{0}=m)\cdot\mathbb{E}[\|\mathbf{x}_{r}(s)-W_{d}\sigma(W_{e}\mathbf{x}_{p}(s))\|^{2}\mid\|\mathbf{x}(s)\|_{0}=m]$ (27)
$=\cdot\sum_{d=1}^{n}M_{d}\left\|w_{r}^{d}-W_{d}\sigma(W_{e}w_{p}^{d})\right\|^{2}+O((1-S)^{2})$ (28)
Appendix B Proof of Theorem 3.2
Proof.
We analyze the approximate loss $\tilde{\mathcal{L}}_{\text{SDL}}(W_{d}^{*},W_{e}^{*})$ by examining the reconstruction error for each ground-truth feature when active in isolation.
Consider feature $d\in[n]$ active in isolation. The encoder output is $W_{e}^{*}w_{p}^{d}=W_{p}^{\top}w_{p}^{d}$ with $k$-th component:
$(W_{e}^{*}w_{p}^{d})_{k}=\langle w_{p}^{k},w_{p}^{d}\rangle$ (29)
Under Assumption 2.4, the unit-norm condition gives $(W_{e}^{*}w_{p}^{d})_{d}=1$. For $k\neq d$, the bounded interference condition implies $\langle w_{p}^{k},w_{p}^{d}\rangle\leq M$, where $M$ is the maximum interference. Combined with the Cauchy-Schwarz inequality, we have:
$\langle w_{p}^{k},w_{p}^{d}\rangle\in[-1,M]\quad\forall k\neq d$ (30)
Define the set of activated neurons for feature $d$:
$\mathcal{A}_{d}=\{k\in[n]:\sigma(W_{e}^{*}w_{p}^{d})_{k}>0\}$ (31)
Since $\sigma(z)_{i}\in\{0,z_{i}\}$ and typical activation functions satisfy $\sigma(z)_{i}>0\iff z_{i}>c$ for some threshold $c\geq 0$, we have:
$\mathcal{A}_{d}=\{k\in[n]:\langle w_{p}^{k},w_{p}^{d}\rangle>c\}$ (32)
Note that $d\in\mathcal{A}_{d}$ since $(W_{e}^{*}w_{p}^{d})_{d}=1>c\geq 0$. For $k\neq d$, neuron $k$ activates only if $\langle w_{p}^{k},w_{p}^{d}\rangle>c\geq 0$, which requires positive interference. Therefore:
$\mathcal{A}_{d}\setminus\{d\}\subseteq\{k\in[n]:0<\langle w_{p}^{k},w_{p}^{d}\rangle\leq M\}$ (33)
Let $K_{d}=|\mathcal{A}_{d}|$ denote the number of activated neurons.
The activation output satisfies $\sigma(W_{e}^{*}w_{p}^{d})_{k}=\langle w_{p}^{k},w_{p}^{d}\rangle$ for $k\in\mathcal{A}_{d}$ and zero otherwise, since $\sigma(z)_{i}\in\{0,z_{i}\}$. The reconstruction is therefore $W_{d}^{*}\sigma(W_{e}^{*}w_{p}^{d})=\sum_{k\in\mathcal{A}_{d}}\langle w_{p}^{k},w_{p}^{d}\rangle w_{r}^{k}$. The reconstruction error for feature $d$ is
$w_{r}^{d}-W_{d}^{*}\sigma(W_{e}^{*}w_{p}^{d})$ $=w_{r}^{d}-\sum_{k\in\mathcal{A}_{d}}\langle w_{p}^{k},w_{p}^{d}\rangle w_{r}^{k}$ (34)
$=w_{r}^{d}-w_{r}^{d}-\sum_{k\in\mathcal{A}_{d}\setminus\{d\}}\langle w_{p}^{k},w_{p}^{d}\rangle w_{r}^{k}$ (35)
$=-\sum_{k\in\mathcal{A}_{d}\setminus\{d\}}\langle w_{p}^{k},w_{p}^{d}\rangle w_{r}^{k}$ (36)
Under the unit-norm assumption, $\|w_{r}^{k}\|=1$ for all $k$. For $k\in\mathcal{A}_{d}\setminus\{d\}$, we have $\langle w_{p}^{k},w_{p}^{d}\rangle>0$ (since negative interference does not activate) and $\langle w_{p}^{k},w_{p}^{d}\rangle\leq M$ by the bounded interference condition. Therefore, using the triangle inequality:
$\left\|w_{r}^{d}-W_{d}^{*}\sigma(W_{e}^{*}w_{p}^{d})\right\|$ $\leq\sum_{k\in\mathcal{A}_{d}\setminus\{d\}}|\langle w_{p}^{k},w_{p}^{d}\rangle|\cdot\|w_{r}^{k}\|$ (37)
$=\sum_{k\in\mathcal{A}_{d}\setminus\{d\}}\langle w_{p}^{k},w_{p}^{d}\rangle\cdot 1\quad\text{(positive terms)}$ (38)
$\leq\sum_{k\in\mathcal{A}_{d}\setminus\{d\}}M$ (39)
$=M(K_{d}-1)$ (40)
Squaring both sides: $\|w_{r}^{d}-W_{d}^{*}\sigma(W_{e}^{*}w_{p}^{d})\|^{2}\leq M^{2}(K_{d}-1)^{2}$. The approximate loss is
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{d}^{*},W_{e}^{*})=\sum_{d=1}^{n}M_{d}\cdot M^{2}(K_{d}-1)^{2}\leq M^{2}\sum_{d=1}^{n}M_{d}(K_{d}-1)^{2}$ (41)
Since $K_{d}\geq 1$ for all $d$ and typically $K_{d}=O(1)$ for most features (as the activation function provides sparsity), we have $(K_{d}-1)^{2}=O(1)$. Under the feature independence and stability assumptions (Assumption 2.4), $\sum_{d=1}^{n}M_{d}=\sum_{d=1}^{n}\Pr(\mathbf{x}(s)=x_{d}(s)e_{d})\cdot\mathbb{E}[x_{d}(s)^{2}|x_{d}(s)>0]=O(1-S)$, since the probability that exactly one feature is active is (1S)(1-S) under extreme sparsity. Therefore:
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{d}^{*},W_{e}^{*})=O(M^{2}(1-S))$ (42)
This completes the proof. ∎
Appendix C Proof of Theorem 3.5
Proof.
We construct a point $(W_{d},W_{e})$ exhibiting the polysemanticity pattern and verify it satisfies all three conditions.
Step 1: Construction of encoder $W_{e}$. Given a realizable polysemanticity pattern $(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$, by Definition LABEL:def:realizable-pattern, there exists an encoder $W_{e}=[(w_{e}^{1})^{\top},\ldots,(w_{e}^{n_{q}})^{\top}]^{\top}$ such that for all $i\in[n_{q}]$:
$\mathcal{F}_{i}=\left\{d\in[n]:\left(\sigma(W_{e}w_{p}^{d})\right)_{i}>0\right\}.$ (43)
Fix this $W_{e}$ for the remainder of the proof. Define the activation vectors:
$z_{d}:=\sigma(W_{e}w_{p}^{d})\in\mathbb{R}^{n_{q}},\quad d=1,\ldots,n.$ (44)
By construction, $(z_{d})_{i}>0$ if and only if $d\in\mathcal{F}_{i}$.
Partition the neurons into active and dead sets:
$\mathcal{A}=\{i\in[n_{q}]:\mathcal{F}_{i}\neq\emptyset\},\quad\mathcal{D}=\{i\in[n_{q}]:\mathcal{F}_{i}=\emptyset\}.$ (45)
Step 2: Gradient analysis with respect to $W_{d}$. The gradient of the loss with respect to the $i$-th decoder column $w_{d}^{i}\in\mathbb{R}^{n_{r}}$ is:
$\nabla_{w_{d}^{i}}\tilde{\mathcal{L}}_{\text{SDL}}=-2\sum_{d=1}^{n}M_{d}(w_{r}^{d}-W_{d}z_{d})(z_{d})_{i}.$ (46)
Case 1: Dead neurons ($i\in\mathcal{D}$). For $i\in\mathcal{D}$, we have $\mathcal{F}_{i}=\emptyset$, which implies $(z_{d})_{i}=0$ for all $d\in[n]$. Therefore:
$\nabla_{w_{d}^{i}}\tilde{\mathcal{L}}_{\text{SDL}}=0\quad\forall w_{d}^{i}.$ (47)
The decoder weights for dead neurons are arbitrary. We set $w_{d}^{i}=0$ for all $i\in\mathcal{D}$.
Case 2: Active neurons ($i\in\mathcal{A}$). For $i\in\mathcal{A}$, we have $\mathcal{F}_{i}\neq\emptyset$. The gradient vanishes when:
$\sum_{d\in\mathcal{F}_{i}}M_{d}w_{r}^{d}(z_{d})_{i}=\sum_{d\in\mathcal{F}_{i}}M_{d}(W_{d}z_{d})(z_{d})_{i}.$ (48)
Expanding the right side and using the fact that $(z_{d})_{j}=0$ for $j\notin\mathcal{F}_{j}$ or $d\notin\mathcal{F}_{j}$:
$\sum_{d\in\mathcal{F}_{i}}M_{d}w_{r}^{d}(z_{d})_{i}=\sum_{j=1}^{n_{q}}w_{d}^{j}\underbrace{\sum_{d\in\mathcal{F}_{i}\cap\mathcal{F}_{j}}M_{d}(z_{d})_{i}(z_{d})_{j}}_{=:A_{ij}}.$ (49)
Define the Gram matrix $A\in\mathbb{R}^{n_{q}\times n_{q}}$:
$A_{ij}=\sum_{d\in\mathcal{F}_{i}\cap\mathcal{F}_{j}}M_{d}(z_{d})_{i}(z_{d})_{j}.$ (50)
And the target matrix $B\in\mathbb{R}^{n_{r}\times n_{q}}$:
$B_{:,i}=\sum_{d\in\mathcal{F}_{i}}M_{d}w_{r}^{d}(z_{d})_{i}.$ (51)
The stationarity condition becomes:
$B=W_{d}A.$ (52)
Step 3: Properties of the Gram matrix $A$.
Claim 3.1: Block structure. The matrix $A$ has a block-diagonal structure with respect to the partition $(\mathcal{A},\mathcal{D})$:
$A=\begin{bmatrix}A_{\mathcal{A}\mathcal{A}}&0\\ 0&0\end{bmatrix}$, (53)
where $A_{\mathcal{A}\mathcal{A}}$ corresponds to rows and columns indexed by $\mathcal{A}$.
Proof of Claim 3.1: For $i\in\mathcal{D}$, we have $\mathcal{F}_{i}=\emptyset$, so $\mathcal{F}_{i}\cap\mathcal{F}_{j}=\emptyset$ for all $j$, giving $A_{ij}=0$. Similarly, $A_{ji}=0$. For $i,j\in\mathcal{D}$, $A_{ij}=0$. $\square$
Claim 3.2: Positive definiteness. The submatrix $A_{\mathcal{A}\mathcal{A}}$ is positive definite.
Proof of Claim 3.2: For any nonzero $v\in\mathbb{R}^{|\mathcal{A}|}$, consider:
$v^{\top}A_{\mathcal{A}\mathcal{A}}v$ $=\sum_{i,j\in\mathcal{A}}v_{i}v_{j}A_{ij}$ (54)
$=\sum_{i,j\in\mathcal{A}}v_{i}v_{j}\sum_{d\in\mathcal{F}_{i}\cap\mathcal{F}_{j}}M_{d}(z_{d})_{i}(z_{d})_{j}$ (55)
$=\sum_{d=1}^{n}M_{d}\left(\sum_{i\in\mathcal{A}:d\in\mathcal{F}_{i}}v_{i}(z_{d})_{i}\right)^{2}.$ (56)
Since $(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ forms a partition of $[n]$, for each $d\in[n]$, there exists at least one $i\in\mathcal{A}$ with $d\in\mathcal{F}_{i}$ and $(z_{d})_{i}>0$. For any nonzero $v$, at least one term in the sum is positive, giving $v^{\top}A_{\mathcal{A}\mathcal{A}}v>0$. $\square$
Step 4: Optimal decoder solution. From the normal equations (52), the optimal decoder for active neurons is:
$W_{d}^{*}[:,\mathcal{A}]=B[:,\mathcal{A}](A_{\mathcal{A}\mathcal{A}})^{-1},$ (57)
where $(A_{\mathcal{A}\mathcal{A}})^{-1}$ exists by Claim 3.2. For dead neurons, we set:
$W_{d}^{*}[:,\mathcal{D}]=0.$ (58)
This gives $\nabla_{W_{d}}\tilde{\mathcal{L}}_{\text{SDL}}(W_{d}^{*},W_{e})=0$, proving condition (2).
Step 5: Hessian analysis. The Hessian with respect to $W_{d}$ has Kronecker structure:
$\nabla^{2}_{W_{d}}\tilde{\mathcal{L}}_{\text{SDL}}=I_{n_{r}}\otimes(2A).$ (59)
For positive semi-definiteness, we need to show $A\succeq 0$. From the block structure (Claim 3.1) and positive definiteness of $A_{\mathcal{A}\mathcal{A}}$ (Claim 3.2):
$A=\begin{bmatrix}A_{\mathcal{A}\mathcal{A}}&0\\ 0&0\end{bmatrix}\succeq 0,$ (60)
which implies $\nabla^{2}_{W_{d}}\tilde{\mathcal{L}}_{\text{SDL}}(W_{d}^{*},W_{e})\succeq 0$, proving condition (3).
Step 6: Non-zero loss for polysemantic patterns. We now prove that $\tilde{\mathcal{L}}_{\text{SDL}}(W_{d}^{*},W_{e})>0$ when the pattern exhibits polysemanticity.
By Theorem LABEL:thm:global-necessary-sufficient, $\tilde{\mathcal{L}}_{\text{SDL}}=0$ if and only if:
$w_{r}^{d}=W_{d}\sigma(W_{e}w_{p}^{d})=W_{d}z_{d}\quad\forall d\in[n].$ (61)
Since the pattern has at least one $|\mathcal{F}_{i}|\geq 2$, fix such an $i\in\mathcal{A}$ and choose $d_{1},d_{2}\in\mathcal{F}_{i}$ with $d_{1}\neq d_{2}$. Both features activate neuron $i$, so $(z_{d_{1}})_{i}>0$ and $(z_{d_{2}})_{i}>0$.
The perfect reconstruction conditions for these features are:
$w_{r}^{d_{1}}$ $=\sum_{j=1}^{n_{q}}w_{d}^{j}(z_{d_{1}})_{j}=\sum_{j\in\mathcal{F}_{j}:d_{1}\in\mathcal{F}_{j}}w_{d}^{j}(z_{d_{1}})_{j},$ (62)
$w_{r}^{d_{2}}$ $=\sum_{j=1}^{n_{q}}w_{d}^{j}(z_{d_{2}})_{j}=\sum_{j\in\mathcal{F}_{j}:d_{2}\in\mathcal{F}_{j}}w_{d}^{j}(z_{d_{2}})_{j}.$ (63)
Both sums include the term $w_{d}^{i}(z_{d_{k}})_{i}$ for $k\in\{1,2\}$. However, the coefficients differ: $(z_{d_{1}})_{i}\neq(z_{d_{2}})_{i}$ in general (since different input patterns).
Key observation: Under Assumption 2.4, $w_{r}^{d_{1}}\not\propto w_{r}^{d_{2}}$ for distinct features. If both (61) equations held, we would need:
$\frac{w_{r}^{d_{1}}}{(z_{d_{1}})_{i}}=w_{d}^{i}=\frac{w_{r}^{d_{2}}}{(z_{d_{2}})_{i}},$ (64)
which requires $w_{r}^{d_{1}}\propto w_{r}^{d_{2}}$, contradicting Assumption 2.4.
Therefore, the optimal decoder $W_{d}^{*}$ cannot satisfy (61) for all $d$, giving:
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{d}^{*},W_{e})=\sum_{d=1}^{n}M_{d}\|w_{r}^{d}-W_{d}^{*}z_{d}\|^{2}>0,$ (65)
proving condition (1).
Conclusion. We have constructed $(W_{d}^{*},W_{e})$ exhibiting the polysemanticity pattern such that all three conditions hold, completing the proof. ∎
Appendix D Proof of Theorem 3.8
Proof.
We show that hierarchical concept structures naturally give rise to polysemanticity patterns with feature absorption, and then apply Theorem 3.5 to establish the existence of local minima.
Consider ground-truth features exhibiting hierarchical structure with parent concept $p$ and sub-concepts $\mathcal{C}=\{c_{1},\ldots,c_{k}\}$ satisfying Definition 3.7. Any partition of $\mathcal{C}$ into non-empty subsets $\mathcal{C}_{1},\mathcal{C}_{2}$ with $\mathcal{C}_{1}\cup\mathcal{C}_{2}=\mathcal{C}$ and $\mathcal{C}_{1}\cap\mathcal{C}_{2}=\emptyset$ induces a polysemanticity pattern as follows.
Define $(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ where:
  • $\mathcal{F}_{i}=\mathcal{C}_{1}$ for some neuron $i\in[n_{q}]$
  • $\mathcal{F}_{j}=\mathcal{C}_{2}$ for some neuron $j\in[n_{q}]$, $j\neq i$
  • $\mathcal{F}_{\ell}=\emptyset$ for all other neurons $\ell\notin\{i,j\}$
This constitutes a polysemanticity pattern since:
  1. 1.
    $(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ forms a partition of $[n]$ restricted to the sub-concepts (assuming other ground-truth features are assigned to other neurons)
  2. 2.
    At least one of $|\mathcal{F}_{i}|\geq 2$ or $|\mathcal{F}_{j}|\geq 2$ holds, satisfying the polysemanticity condition in Definition 3.4
  3. 3.
    The pattern exhibits feature absorption: neuron $i$ captures a subset of sub-concepts (absorbed feature) while neuron $j$ captures the remaining sub-concepts (main line interpretation)
If this polysemanticity pattern is realizable (Definition 3.4), then by Theorem 3.5, there exists a configuration $(W_{d},W_{e})$ that is a local minimum of $\tilde{\mathcal{L}}_{\text{SDL}}$ exhibiting this pattern. This completes the proof. ∎
Appendix E The Linear Representation Bench
Figure 8: Visualization of the Linear Representation Bench. The figure illustrates a $D=3$ dimensional representation space generated using $N=4$ feature directions ($\mathbf{W}_{p}$).
Appendix F Limitations and Future Works
While our theoretical framework provides valuable insights into SDL methods, several limitations warrant discussion:
  • Assumption Violations. Our analysis relies on Assumptions 2.3 and 2.4, which may not hold perfectly in realworld neural networks.
  • Extreme Sparsity. Some of our theoretical results are limited to the extreme sparsity ($S\to 1$) assumption, which may not hold in realworld neural networks.
  • Convergence Guarantees. We characterize the global and local minima of the optimization landscape. Future works can provide convergence guarantees or sample complexity bounds following our framework.
  • Classical SDL Methods. Established dictionary learning algorithms have demonstrated success in signal processing but remain unexplored for mechanistic interpretability, representing promising future directions.